An LED operates at 1.5 V and 5 mA in forward bias. Assuming an 80% external
efficiency of the LED, how many photons are emitted per second?
1
Expert's answer
2020-09-28T08:10:45-0400

"P_{in} = η_{ext}\frac{i}{e}hf"

Number of photons "= \frac{P_{in}}{hf} = η_{ext}\frac{i}{e}"

Number of photons "= 0.8 \times \frac{5 \times 10^{-3}}{1.6 \times 10^{-19}}"

Number of photons "= 2.5 \times 10^{16}"